What is the value of a preferred stock when the dividend rate is 16 percent on a $125 par value?

(Preferred stock valuation) What is the value of a preferred stock when the dividend rate is 16 percent on a $125 par value? The appropriate discount rate for a stock of this risk level is 13 percent. 

The value of the preferred stock is $ _______ . (Round to the nearest cent.) 


(Preferred stock valuation) The preferred stock of Gandt Corporation pays a $3.75 dividend. What is the value of the stock if your required return is 14 percent? 

The value of the preferred stock is $_______  per share. (Round to the nearest cent.)
Answer #1

a)

Preferred Stock Dividend = Face value * Dividend rate = $125*16% =$20

Value of Preferred stock = Preferred stock dividend / Discount rate

Value of Preferred stock = $20/0.13 = $154

b)

Value of Preferred stock = $3.75/0.14 =$27